OpenStudy (anonymous):

x-4((x^2+5)^1/2)=0 solve for x

OpenStudy (anonymous):

\[x-4 \sqrt{5+x^2}=0 \]\[x^2\text{=}\left(4 \sqrt{5+x^2}\right)^2\]\[x^2=80+16 x^2\]\[x=\frac{4 i}{\sqrt{3}} \]

OpenStudy (anonymous):

plus or minus

OpenStudy (anonymous):

Mathematica says the answer is +(4*i)/radical 3